If A = (5√2+6√3)(4√2-5√3) & B = (7√2+5√3)(8√2-5√3) , then find A + B.

Grade:9

1 Answers

Vikas TU
4 years ago
Dear student 
A = (5√2+6√3)(4√2-5√3) & B = (7√2+5√3)(8√2-5√3) 
A = 40 – 25√6 +24√6 -90 
=  -50 – √6 
B = 112 – 35√6 + 40√6 -75 
= 37 + 5 √6 
A + B = -13 + 4√6 
Hope this helps 
Good Luck